NAOJ GW Elog Logbook 3.2

LC 1 | LC 2 | |
P | 0.53 | 0.55 |
I | 0.34 | 0.32 |
D | 0.21 | 0.27 |


A parameter for a phase shifted voltage waveform between the two LCs has also been added.
The VI is completely ready for use. The speed of VI (with all data saving) is 40Hz.

Measured Fast axis orientation of second LC
Issue solved:
1. see elog 3242 for LC 1 fast axis. It was continuous (done in the PP(Parallel Polarizer) configuration)
2. see elog 3254 for LC 2 fast axis it was discontinuous.
3. We did LC 2 measurements and found Fig 1 and Fig 2 using the CP(Cross Polarizer) technique. But we used the cos function to fit it and other mistakes that are addressed below. We see that even if we moved our LC(by -11.62deg ) it seems that the LC didn't rotate at all.
Mistakes we made
1. When using the fit equation the bounds of phase were given from -4 to 4. I have changed it from 0 to 2*pi.
2. We should use the sine function to fit cross-polarizer data and cos for Parallel polarizer. This arises from the fact that we use power to understand the position of the fast axis. So if you observe data from these two techniques like power vs. rotation. It will be shifted.
3. Now, let's compare Fig 1 and Fig 3 where CP data was fit using right parameters.
4. we also need to consider the initial angle where we start measurement both when observing power vs rotation and analyzing fast axis orientation. When I incorporated this we could exactly see how much were we rotating the fast axis or if we rotated in the wrong direction
5. So when we first did a measurement with CP (shown in Fig 1) (we had a lot of data points because we did it with a 10-degree resolution). But because we were using the wrong function to fit no matter how many times we did the measurement our data looked strange
6. When we shifted to PP after 4 sets of measurements, we did measure with less resolution and so it looked strange.
7. We can see from Fig 3 that we are supposed to move LC by 11.43 degree. so I moved LC and measured again to obtain Fig 4, Fig 5 and Fig 6.
8. Just to be sure that our new analysis is correct you can compare Fig 7 from elog 3242 for LC 1's fast axis.
9. As a matter of fact, it's better to use CP for such measurements to see small changes too (because of small retardance range of LC at high voltage). See Fig 8 and Fig 9 for measurement done using PP. Although these measurements entailed the corrections during analysis, the data seemed strange enough to be considered.
Miscellanous:
1. some angles were omitted during analysis (likely because the LC cable caused disruption in beam and hence strange data). Also, this won't matter because we won't use these omitted orientations during characterization. Preferably the LC will be fixed along the fast axis
2. Moving on, the issue of temperature rising in LC was completely unrelated to the box or laser. It was due to us using tape to fix the wire so that it doesn't move and block the beam. This tension caused an elevated temperature of 26 deg even when the laser was off, the temperature controller was off and the box was lifted. The temperature was relieved as soon as I removed the tape!
for PP folder_name = r'C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\LC2_calibration data\fast axis 8'
for CP folder_name = r'C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\LC2_calibration data\fast axis 9'

Summary
When we were working with Yuhang in April, we noticed that the CC loop was very glitchy. I attempted to measure the CC PLL phase noise but got no signal.
Details
I attempted to measure PLL phase noise using the same method I did previously. To recap:
- The PLL outputs a sine wave
- If this is mixed with some local oscillator at the same frequency, then the output of the mixing will just be zero signal + phase noise
- The procedure is to put [name] PLL MON into RF and a local oscillator directly from the DDS3 into LO (last time I used 7.6 dBm).
- DDS3 outputs: Ch0 - CC LO 21 MHz, Ch1 - CC2, Ch2 - CC1, Ch3 - PPol LO 35 MHz - note that Ch1 and Ch3 are reversed compared to wiki/theses.
- Offsetting the PLL by 100 Hz should give a signal visible on the oscilloscope at about 0.08 V
I could see that the LO was generated from the DDS board (Ch3 - 50 MHz, 9 dBm), and that the CC PLL could lock. However, no signal came out of the mixer - the spectrum analyser just showed the same as when unplugged, and putting the signal into the oscilloscope gave nothing when the RF and LO were offset by 100 Hz.
Next time I will check the mixer and cables

The measurement was repeated. First, I moved the LC to the new position of 337.78 degrees. Then I repeated the measurement with the angle of rotation incremented by 30 degrees.
The issue with fast axis orientation continues. see Fig 1 and 2. It seems that there is a discontinuity at some points. This has been observed in almost all the 7 times that we did measurements.
PS:
The data is stored in foldername= 'C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\LC2_calibration data\fast axis 7'

[Shalika, Marc]
This measurement aims to find the fast axis orientation of our second LC so that we can align it correspondingly for future measurements.
1. The fast axis orientation was measured using the method mentioned in elog 3242.
2. Our setup is inside a box and so (sometimes the laser hits the box or because of lack of air circulation) the temperature gets unstable and affects the LC. So we decide to use LC at 30 degrees. This helps us avoid any temperature fluctuations.
3. The LC was rotated from 0 to 360 deg with increments of 45 degrees.
4. To understand our fast axis position our data (where power is observed for the LC rotation) was fitted properly to the equations. (see Fig 1) . This fit is then used to get fast axis orientation.
5. The axis was found to be at -22.3 degrees (see Fig 2) and so the LC has been rotated to this position. The LC position is now 337.8 degrees.
6. The measurement will be taken again to avoid any deviations from the best position, with 10 degree increments.
[we have tried taking this measurement 6 times as of now. We are a bit baffled by the mysterious nature of fast axis discontinuity. We will try to do this measurement with more points. ]
PS:
The data is stored in foldername= 'C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\LC2_calibration data\fast axis 6'
The measurement was repeated. First, I moved the LC to the new position of 337.78 degrees. Then I repeated the measurement with the angle of rotation incremented by 30 degrees.
The issue with fast axis orientation continues. see Fig 1 and 2. It seems that there is a discontinuity at some points. This has been observed in almost all the 7 times that we did measurements.
PS:
The data is stored in foldername= 'C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\LC2_calibration data\fast axis 7'

To measure the fast axis orientation we shuold have the 2 polarizers parallels and not crossed.

[Shalika, Marc]
1. The extinction ratio of our second LC was measured using the method mentioned in elog 3241. The power at transmission was observed to obtain the maximum and minimum power for every voltage by correspondingly rotating the LC.
2. The error bars are coming from the power fluctuations we measured.
3. The extinction ratio observed was at about 1.0332+/-0.0013 for all voltages. [see Fig 1]
4. Also the setup was moved to accomodate the new vertical slider and box because we have overhead table too. [see Fig 2]
[We also measured the fast axis orientation using the method in elog 3242 but the results were a bit confusing so its work in progress]

[Marc, Yuhang {remote)]
We checked the reason why healthcheck failed.
The test.sh code calls a .xml file for each coil of every suspension. All these files are located in the check_after_earthquake folder.
These .xml files are used to set up the measurement, from excitation to measurement.
However, no measurement nor excitation channels were specified aswell as reference measurements.
Looking at various folders, we found the correct files.
They were restored to the proper folder and we confirmed that the suspension health-check is recovered.

[Hirata, Marc, Shalika]
In order to ease the installation of optics for birefringence measurement we added a vertical slidder to the box.
We use an unused rack slider fixed to an unused part of the old scattering frame.
To stop the box at a higher position it is good to use 2 keys for safety.

[Marc, Michael]
To prepare the shipping of the AOM function generator back to APC, we installed our new one (WF1968).
We used same settings as before (continuous mode, freqeuncy = 109.036035615 MHz, 1.124Vpk=5dBm).
We had quite low green power after the IRIS after the AOM.
Power budget was :
SHG reflection ~270mW
before AOM ~41.2mW
after AOM ~37.6 mW
after iris ~5.7mW
before FC ~5.26 mW.
We tuned a bit the steering mirror between the AOM and iris. Especially, there seemed to be a little clipping in the vertical direction but it did not help to recover the usual power.
We checked the function generator output on spectrum analyzer and found a -3.28 dBm amplitude. Therefore, we increased the amplitude to recover the usual 5dBM : 2.95VPk gives 5.1 dBm.
In this configuration we recovered the usual AOM transmission.
We had to realign the last steering mirror on the bench to recover the PR targets.
The beam was then really far of the BS target. We tried to move PR mirror bu it did not seem to move. The health check has some error message ('No measurement channel defined') that we need to investigate.

The overhead sheld (ATS6) was delivered and assembled by Japan Laser people.
It is possible to move the hanging shelf and the cable management tray by unscrewing the bolts.
Be careful that there are other bolts inside the frame that need to be moved as well.
There are also extra bolts and smaller tube for the hanging shelf in a plastic bag taped to the optical table foot.

[Marc, Takahashi, Yoshizumi]
We removed the failling tape and replaced them with plastic ones.
The pre-clean room and clean room wall are now repaired.

[Marc, Munetake, Shalika]
We reconfigured the PCI for birefringence measurement.
First, we put back all required cables with the 6dB attenuators at the output of the PSD sum.
We tuned the laser power to 1.6mW (laser current = 1A and HWP rotated to 348deg.
We installed the razor blades and tuned the vertical and horizontal angle of incidence to about 0.006deg.
We realigned the readout part.
We tuned the QWP and HWP to minimize p pol in transmission with HWP = 351.2 deg.
We took 10mn of s and p polarization and got the attached calibration coefficients.

[Marc, Shalika]
Following the several months with the too large scattering box, the wall of the PCI clean room are quite damaged...
We should definitely replace them.
We removed the box and several of the optics used for the scattering measurement.
Because of the damaged wall, the 'clean room' was quite dirty and we spent some time cleaning it up.
We install a new shelf in front of the PCI clean room entrance to store the scattering components
[Marc, Takahashi, Yoshizumi]
We removed the failling tape and replaced them with plastic ones.
The pre-clean room and clean room wall are now repaired.

I started the SIP "N-S P8" near the south end. Applied voltag and current were changed as follows.
[Just after starting]
N-S P8 | |
Voltage [V] | 5910 |
Current [mA] | 0.8 |
[After 3 hours]
N-S P8 | |
Voltage [V] | 5910 |
Current [mA] | 0.4 |


[Marc, Shalika]
When we measured the polarization state with the polarization camera directly after the LC we could see that the azimuth angle seems to change a lot around the half-wave retardation voltage.
We tried to measure the fast axis direction by minimizing the transmitted power of the LC after a polarizer but the results were quite strange and not so consistent when repeated..
One possible explanation is that the power minimization was not precise enough to do by hand.
We decided to follow the procedure of [1] where we installed the output polarizer to be in cross-polarizer configuration and measured the transmitted power while appling a sawtooth voltage and rotating the LC from 0 to 360 deg with 10deg increment.
The resulting function is fitted for a given voltage by a sum of cosine with different orders, all as a function of (x-x0) where x is the rotation angle of the LC and x0 a possibly voltage-dependent offset that should correspond to the LC fast axis rotation as a function of the applied voltage.
By repeating this for every voltage we can get the attached figure where we found that the fast axis orientation is almost voltage independent (within 1deg).
[1] : Measurements of linear diattenuation and linear retardance spectra with a rotating sample spectropolarimeter David B. Chenault and Russell A. Chipman
To measure the fast axis orientation we shuold have the 2 polarizers parallels and not crossed.

[Marc, Shalika]
To measure the extinction ratio of our LC, we removed the polarizer and rotated the LC to measure the maximum and minimum of the transmitted power.
To be more precise, the transmitted power is normalized by the input power.
We repeated this measurement for several voltages applied to the LC as in the attached figure.
The errorbars are coming from the power fluctuations we measured.
The extinction ratio seems quite constant at all voltages at about 1.009+/-0.002